# Well-posedness of a bulk-surface convective Cahn–Hilliard system with dynamic boundary conditions

### v2

@@ 1   Introduction
The function $\phi:Q\rightarrow\mathbb{R}$ is the bulk phase-field. It is an order parameter, which represents the distribution of the two immiscible materials within the domain $\Omega$. Similarly, the surface phase-field $\psi:\Sigma\rightarrow\mathbb{R}$ represents the distribution of two materials on the surface. The functions $F$ and $G$ are double-well potentials, which lead to the effect that $\phi(t)$ and $\psi(t)$ will attain values close to ±$\pm 1$ in most parts of $\Omega$ and $\Gamma$ (which correspond to the pure phases of the materials). In the remaining intermediate regions, which are called diffuse interfaces, $\phi$ and $\psi$ are expected to exhibit a continuous transition between the values $-1$ and $1$. The functions $\mu:Q\rightarrow\mathbb{R}$ and $\theta:\Sigma\rightarrow\mathbb{R}$ represent the bulk chemical potential and the surface chemical potential, respectively. The non-negative, scalar functions $m_{\Omega}(\phi)$ and $m_{\Gamma}(\psi)$ are called mobilities. They depend on the phase-fields and describe where diffusion processes occur. Moreover, the functions $\boldsymbol{v}:Q\rightarrow\mathbb{R}^{d}$ and $\boldsymbol{w}:\Sigma\rightarrow\mathbb{R}^{d}$ are prescribed velocity fields that correspond to the flow of the two materials in the bulk and on the surface, respectively. In this paper, we always assume that $\boldsymbol{v}\cdot\mathbf{n}=0$ on $\Sigma$. We point out that in many cases (e.g., in the model derived in [23]), $\boldsymbol{v}$ and $\boldsymbol{w}$ are related by the condition $\boldsymbol{w}=\boldsymbol{v}|_{\Sigma}$ on $\Sigma$. However, as this relation will not have any impact on our mathematical analysis, we consider $\boldsymbol{v}$ and $\boldsymbol{w}$ as independent functions to keep our model as general as possible. Furthermore, $\epsilon>0$ is a parameter related to the thickness of the diffuse interface in the bulk, whereas $\epsilon_{\Gamma}>0$ corresponds to the width of the diffuse interface on the boundary. The parameter $\kappa>0$ acts as a weight for the surface Dirichlet energy $\psi\mapsto\int_{\Gamma}\left|\nabla_{\Gamma}\psi\right|^{2}\;\mathrm{d}\Gamma$, which has a smoothing effect on the phase separation at the boundary.
@@ 2.1   Notation
    For any bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\in\mathbb{N}$) with Lipschitz boundary $\Gamma:=\partial\Omega$, we introduce the space
    $\mathbf{L}^{3}_{\textnormal{div}}(\Omega)\coloneqq\big{\{}\boldsymbol{v}\in\mathbf{L}^{3}(\Omega):\textnormal{div}\;\boldsymbol{v}=0\ \text{in\leavevmode\nobreak\ }\Omega,\ \boldsymbol{v}\cdot\mathbf{n}=0\ \text{on\leavevmode\nobreak\ }\Gamma\big{\}}.$
@@ 3   Main results
Suppose that assumptions (A1)(A4) hold. Let $K,L\in[0,\infty]$, let $(\phi_{0},\psi_{0})\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data and let $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$ be given velocity fields. The quadruplet $(\phi,\psi,\mu,\theta)$ is called a weak solution of the system (3.1) on $[0,T]$ if the following properties hold:
@@
(Existence of weak solutions of (3.1)) Suppose that the assumptions (A1)(A4) hold. Let $K,L\in[0,\infty]$, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data, and let $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$ be given velocity fields. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. Then, there exists a weak solution $(\phi,\psi,\mu,\theta)$ of the system (3.1) in the sense of Definition 3.1, which has the additional regularity
@@ 4   Existence of weak solutions in the case $K,L\in(0,\infty)$
Suppose that the assumptions (A1)(A4) hold. Let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}$ be arbitrary initial data, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$, $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$ and $K,L\in(0,\infty)$. Then there exists a weak solution of the system (3.1) in the sense of Definition 3.1 which further satisfies $\left(\mu,\theta\right)\in L^{4}(0,T;\mathcal{L}^{2})$.
@@ 5.1   The limit $K\rightarrow 0$ and the existence of a weak solution if $(K,L)\in\{0\}\times(0,\infty)$
(Asymptotic limit $K\rightarrow 0$) Suppose that the assumptions (A1)(A5) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{D}_{\alpha}$ be arbitrary initial data, $L\in(0,\infty)$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$. For any $K\in(0,\infty)$, let $(\phi_{K},\psi_{K},\mu_{K},\theta_{K})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi_{\ast},\psi_{\ast},\mu_{\ast},\theta_{\ast})$ with $\phi_{\ast}=\alpha\psi_{\ast}$ a.e. on $\Sigma$ such that
@@
(Asymptotic limit $K\rightarrow\infty$) Suppose that the assumptions (A1)(A5) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}$ be arbitrary initial data, $L\in(0,\infty)$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$. For any $K\in(0,\infty)$, let $(\phi_{K},\psi_{K},\mu_{K},\theta_{K})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi^{\ast},\psi^{\ast},\mu^{\ast},\theta^{\ast})$ such that
@@ 5.3   The limit $L\rightarrow 0$ and the existence of a weak solution if $(K,L)\in[0,\infty]\times\{0\}$
(Asymptotic limit $L\rightarrow 0$) Suppose that the assumptions (A1)(A4) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data for $K\in[0,\infty]$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. For any $L\in(0,\infty)$, let $(\phi_{L},\psi_{L},\mu_{L},\theta_{L})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi_{\ast},\psi_{\ast},\mu_{\ast},\theta_{\ast})$ with $\mu_{\ast}=\beta\theta_{\ast}$ a.e. on $\Sigma$ such that
@@
(Asymptotic limit $L\rightarrow\infty$) Suppose that the assumptions (A1)(A4) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data for $K\in[0,\infty]$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. For any $L\in(0,\infty)$, let $(\phi^{L},\psi^{L},\mu^{L},\theta^{L})$ denote a weak solution to the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then there exists a quadruplet $(\phi^{\ast},\psi^{\ast},\mu^{\ast},\theta^{\ast})$ such that
@@ 6.1   Higher spatial regularity for the phase-fields
(Higher regularity) Suppose that the assumptions (A1)(A6) hold, and let $K,L\in[0,\infty]$, $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}(\Gamma))$. Suppose that the domain $\Omega$ is of class $C^{k}$ for $k\in\{2,3\}$, and in the case $d=3$, we further assume that (A4) holds with $p\leq 4$. Then, if a weak solution $(\phi,\psi,\mu,\theta)$ of the system (3.1) in the sense of Definition 3.1 exists, it additionally satisfies

### v3

@@ 1   Introduction
The function $\phi:Q\rightarrow\mathbb{R}$ is the bulk phase-field. It is an order parameter, which represents the distribution of the two immiscible materials within the domain $\Omega$. Similarly, the surface phase-field $\psi:\Sigma\rightarrow\mathbb{R}$ represents the distribution of two materials on the surface. The functions $F$ and $G$ are double-well potentials, which lead to the effect that $\phi(t)$ and $\psi(t)$ will attain values close to ±$\pm 1$ in most parts of $\Omega$ and $\Gamma$ (which correspond to the pure phases of the materials). In the remaining intermediate regions, which are called diffuse interfaces, $\phi$ and $\psi$ are expected to exhibit a continuous transition between the values $-1$ and $1$. The functions $\mu:Q\rightarrow\mathbb{R}$ and $\theta:\Sigma\rightarrow\mathbb{R}$ represent the bulk chemical potential and the surface chemical potential, respectively. The non-negative, scalar functions $m_{\Omega}(\phi)$ and $m_{\Gamma}(\psi)$ are called mobilities. They depend on the phase-fields and describe where diffusion processes occur. Moreover, the functions $\boldsymbol{v}:Q\rightarrow\mathbb{R}^{d}$ and $\boldsymbol{w}:\Sigma\rightarrow\mathbb{R}^{d}$ are prescribed velocity fields that correspond to the flow of the two materials in the bulk and on the surface, respectively. In this paper, we always assume that $\boldsymbol{v}\cdot\mathbf{n}=0$ and $\boldsymbol{w}\cdot\mathbf{n}=0$ on $\Sigma$. We point out that in many cases (e.g., in the model derived in [23]), $\boldsymbol{v}$ and $\boldsymbol{w}$ are related by the condition $\boldsymbol{w}=\boldsymbol{v}|_{\Sigma}$ on $\Sigma$. However, as this relation will not have any impact on our mathematical analysis, we consider $\boldsymbol{v}$ and $\boldsymbol{w}$ as independent functions to keep our model as general as possible. Furthermore, $\epsilon>0$ is a parameter related to the thickness of the diffuse interface in the bulk, whereas $\epsilon_{\Gamma}>0$ corresponds to the width of the diffuse interface on the boundary. The parameter $\kappa>0$ acts as a weight for the surface Dirichlet energy $\psi\mapsto\int_{\Gamma}\left|\nabla_{\Gamma}\psi\right|^{2}\;\mathrm{d}\Gamma$, which has a smoothing effect on the phase separation at the boundary.
@@ 2.1   Notation
    For any bounded domain $\Omega\subset\mathbb{R}^{d}$ ($d\in\mathbb{N}$) with Lipschitz boundary $\Gamma:=\partial\Omega$, we introduce the spaces
    $\mathbf{L}^{3}_{\textnormal{div}}(\Omega)$$\coloneqq\big{\{}\boldsymbol{v}\in\mathbf{L}^{3}(\Omega):\textnormal{div}\;\boldsymbol{v}=0\ \text{in\leavevmode\nobreak\ }\Omega,\ \boldsymbol{v}\cdot\mathbf{n}=0\ \text{on\leavevmode\nobreak\ }\Gamma\big{\}},$
    $\mathbf{L}^{3}_{\tau}(\Gamma)$$\coloneqq\big{\{}\boldsymbol{w}\in\mathbf{L}^{3}_{\tau}(\Gamma):\boldsymbol{w}\cdot\mathbf{n}=0\ \text{on\leavevmode\nobreak\ }\Gamma\big{\}}.$
@@ 3   Main results
Suppose that assumptions (A1)(A4) hold. Let $K,L\in[0,\infty]$, let $(\phi_{0},\psi_{0})\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data and let $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$ be given velocity fields. The quadruplet $(\phi,\psi,\mu,\theta)$ is called a weak solution of the system (3.1) on $[0,T]$ if the following properties hold:
@@
(Existence of weak solutions of (3.1)) Suppose that the assumptions (A1)(A4) hold. Let $K,L\in[0,\infty]$, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data, and let $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$ be given velocity fields. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. Then, there exists a weak solution $(\phi,\psi,\mu,\theta)$ of the system (3.1) in the sense of Definition 3.1, which has the additional regularity
@@ 4   Existence of weak solutions in the case $K,L\in(0,\infty)$
Suppose that the assumptions (A1)(A4) hold. Let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}$ be arbitrary initial data, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$, $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$ and $K,L\in(0,\infty)$. Then there exists a weak solution of the system (3.1) in the sense of Definition 3.1 which further satisfies $\left(\mu,\theta\right)\in L^{4}(0,T;\mathcal{L}^{2})$.
@@ 5.1   The limit $K\rightarrow 0$ and the existence of a weak solution if $(K,L)\in\{0\}\times(0,\infty)$
(Asymptotic limit $K\rightarrow 0$) Suppose that the assumptions (A1)(A5) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{D}_{\alpha}$ be arbitrary initial data, $L\in(0,\infty)$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$. For any $K\in(0,\infty)$, let $(\phi_{K},\psi_{K},\mu_{K},\theta_{K})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi_{\ast},\psi_{\ast},\mu_{\ast},\theta_{\ast})$ with $\phi_{\ast}=\alpha\psi_{\ast}$ a.e. on $\Sigma$ such that
@@
(Asymptotic limit $K\rightarrow\infty$) Suppose that the assumptions (A1)(A5) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}$ be arbitrary initial data, $L\in(0,\infty)$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$. For any $K\in(0,\infty)$, let $(\phi_{K},\psi_{K},\mu_{K},\theta_{K})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi^{\ast},\psi^{\ast},\mu^{\ast},\theta^{\ast})$ such that
@@ 5.3   The limit $L\rightarrow 0$ and the existence of a weak solution if $(K,L)\in[0,\infty]\times\{0\}$
(Asymptotic limit $L\rightarrow 0$) Suppose that the assumptions (A1)(A4) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data for $K\in[0,\infty]$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. For any $L\in(0,\infty)$, let $(\phi_{L},\psi_{L},\mu_{L},\theta_{L})$ denote a weak solution of the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then, there exists a quadruplet $(\phi_{\ast},\psi_{\ast},\mu_{\ast},\theta_{\ast})$ with $\mu_{\ast}=\beta\theta_{\ast}$ a.e. on $\Sigma$ such that
@@
(Asymptotic limit $L\rightarrow\infty$) Suppose that the assumptions (A1)(A4) hold, let $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$ be arbitrary initial data for $K\in[0,\infty]$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$. In the case $K\in\{0,\infty\}$, we further assume that (A5) holds. For any $L\in(0,\infty)$, let $(\phi^{L},\psi^{L},\mu^{L},\theta^{L})$ denote a weak solution to the system (3.1) in the sense of Definition 3.1 with initial data $\left(\phi_{0},\psi_{0}\right)$. Then there exists a quadruplet $(\phi^{\ast},\psi^{\ast},\mu^{\ast},\theta^{\ast})$ such that
@@ 6.1   Higher spatial regularity for the phase-fields
(Higher regularity) Suppose that the assumptions (A1)(A6) hold, and let $K,L\in[0,\infty]$, $\left(\phi_{0},\psi_{0}\right)\in\mathcal{H}^{1}_{K,\alpha}$, $\boldsymbol{v}\in L^{2}(0,T;\mathbf{L}_{\textnormal{div}}^{3}(\Omega))$ and $\boldsymbol{w}\in L^{2}(0,T;\mathbf{L}^{3}_{\tau}(\Gamma))$. Suppose that the domain $\Omega$ is of class $C^{k}$ for $k\in\{2,3\}$, and in the case $d=3$, we further assume that (A4) holds with $p\leq 4$. Then, if a weak solution $(\phi,\psi,\mu,\theta)$ of the system (3.1) in the sense of Definition 3.1 exists, it additionally satisfies